OpenStudy (anonymous):

Write each equation in standard form. y+6=(x+4) y+3=2/5(x+9)

OpenStudy (anonymous):

i believe that y+6=(x+4) is x+y=-2

OpenStudy (anonymous):

and y+3=2/5(x+9) is 2x - 5y = -3

OpenStudy (anonymous):

do u understand how to write them in standard form? standard form: Ax+By=C

OpenStudy (anonymous):

Can you show me how you did them?

OpenStudy (anonymous):

Or explain?

OpenStudy (anonymous):

yes, if u want them to be in standard form, its always going to end up as Ax+By=C. the first one y+6=(x+4), your going to subtract 6 from both sides, so u subtract 6 by 6 and 4. and u get y=x-2. then u subtract x from both sides and u get -x+y=-2, and that's your answer. im sorry i just noticed that above, i didn't put a negative sign before the x.